OpenStudy (anonymous):

how do i compute the exact value of sin(-10pi/3)

OpenStudy (jamesj):

remember that sin(x) = sin(2pi + x). That will help you.

OpenStudy (anonymous):

Exact value means numerical approximation?

OpenStudy (jamesj):

exact value means exact, closed form value.

OpenStudy (jamesj):

e.g., sin(pi/3) = sqrt(3)/2

OpenStudy (jamesj):

that's the exact value of sin(pi/3)

OpenStudy (jamesj):

Anyway, use sin(x) = sin(2pi + x)

OpenStudy (anonymous):

the answer is in fraction form for example root3/3 how do i get it to that?

OpenStudy (anonymous):

draw unit circle

OpenStudy (anonymous):

ok i have one next to me

OpenStudy (anonymous):

\[\sin(-\frac{10\pi}{3})=\sin(\frac{-12\pi}{3}+\frac{2\pi}{3})=\sin(-4\pi+\frac{2\pi}{3})=\sin(\frac{2\pi}{3})\] \[\sin\left(\frac{2\pi}{3}\right)=\left(\frac{\sqrt{3}}{2}\right)\]